### v2

# DPO: Dual-Perturbation Optimization for Test-time Adaptation in 3D Object Detection

Abstract.
LiDAR-based 3D object detection has seen impressive advances in recent times. However, deploying trained 3D detectors in the real world often yields unsatisfactory performance when the distribution of the test data significantly deviates from the training data due to different weather conditions, object sizes, etc. A key factor in this performance degradation is the diminished generalizability of pre-trained models, which creates a sharp loss landscape during training. Such sharpness, when encountered during testing, can precipitate significant performance declines, even with minor data variations. To address the aforementioned challenges, we propose dual-perturbation optimization (DPO) for Test-time Adaptation in 3D Object Detection (TTA-3OD). We minimize the sharpness to cultivate a flat loss landscape to ensure model resiliency to minor data variations, thereby enhancing the generalization of the adaptation process. To fully capture the inherent variability of the test point clouds, we further introduce adversarial perturbation to the input BEV features to better simulate the noisy test environment. As the dual perturbation strategy relies on trustworthy supervision signals, we utilize a reliable Hungarian matcher to filter out pseudo-labels sensitive to perturbations. Additionally, we introduce early Hungarian cutoff to avoid error accumulation from incorrect pseudo-labels by halting the adaptation process. Extensive experiments across three types of transfer tasks demonstrate that the proposed DPO significantly surpasses previous state-of-the-art approaches, specifically on Waymo $\rightarrow$ KITTI, outperforming the most competitive baseline by 57.72% in $\text{AP}_{\text{3D}}$ and reaching 91% of the fully supervised upper bound. Our code is available at https://github.com/Jo-wang/DPO.
Test-time Adaptation, 3D Object Detection.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28–November 1, 2024; Melbourne, VIC, Australia.booktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28–November 1, 2024, Melbourne, VIC, Australiaisbn: 979-8-4007-0686-8/24/10doi: 10.1145/XXXXXX.XXXXXXjournalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681040isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Scene understandingccs: Computing methodologies Online learning settingsccs: Computing methodologies Object detectionccs: Computing methodologies Transfer learning
1. Introduction
LiDAR-based 3D object detection has gained significant attention with the rapid advancements in autonomous driving (Li et al., 2021; Deng et al., 2021; Luo et al., 2023a; Qian et al., 2022; Luo et al., 2023b; Ju et al., 2022) and robotics (Montes et al., 2020; Zhou et al., 2022), where mainstream 3D detectors are developed to interpret pure point clouds or fuse multimodal knowledge, commonly incorporating camera images (Wang et al., 2023a; Chen et al., 2021b; Liu et al., 2023). However, deploying either point clouds-based or multimodal 3D detection models in real-world scenarios often leads to performance degradation due to distribution shifts between the training data and the encountered real-world data. For instance, a 3D detector trained on the nuScenes dataset (Caesar et al., 2020) might suffer a performance drop when applied to the KITTI dataset (Geiger et al., 2012) due to variations in object sizes and the number of beams. This is known as cross-dataset shift. Additionally, the shift can arise from real-world disturbances, termed as corruption-based shift (Hahner et al., 2022; Kong et al., 2023; Dong et al., 2023), which includes challenges like diverse weather conditions and sensor malfunctions. Moreover, multiple factors are likely to be concurrent, for instance, deploying a 3D detector in a different city while suffering severe snow. This scenario is termed as composite domain shift.
Domain adaption has been discovered (Wang et al., 2022b, 2023b; Luo et al., 2023c; Chen et al., 2021a) to mitigate the performance gap brought by various domain shifts. In 3D object detection, this involves aligning features between the labeled training data and the shifted test data to learn a domain-invariant representation (Luo et al., 2021; Zhang et al., 2021; Zeng et al., 2021) or conducting self-training with the aid of selected pseudo-labels (Chen et al., 2023; Yang et al., 2021, 2022; Peng et al., 2023; Li et al., 2023a). However, these approaches necessitate extensive training over multiple epochs on both training and test sets, rendering them impractical for adaptation to the streaming data. Moreover, the exposure of the training data can significantly compromise its privacy, especially when it contains sensitive user information (e.g., user vehicle trajectories and individuals).
To bridge the performance gap induced by domain shifts, while safeguarding the training data privacy and enabling swift adaptation, test-time adaptation (TTA) emerges as an ideal solution. Prior research on TTA typically adapts a source pre-trained model to the unlabeled test data, either through updating a selected subset of parameters (e.g., BatchNorm layers) (Wang et al., 2021; Niu et al., 2023; Sun et al., 2023), or employing the mean-teacher model (Wang et al., 2022a, 2023d; Yuan et al., 2023; Tomar et al., 2023a) within a single epoch. However, these TTA works currently applied in image classification are inadequate for addressing the dual demands (i.e., object localization and classification) for supervision signals inherent in detection tasks. Within this context, MemCLR (VS et al., 2023) stands out by refining the Region of Interest (RoI) features of detected objects through a transformer-based memory module for 2D object detection. Nevertheless, the stored target representations derived from the source pre-trained model cause performance degradation due to distribution shifts. These limitations pose significant challenges in utilizing previous TTA techniques for 3D object detection.
To tackle these challenges, our goal is to devise an effective strategy for adapting the 3D detection model to various data shifts. We observe a common performance decline when the model encounters unfamiliar scenes. This degradation primarily occurs as the model tends to converge to sharp minima in the loss landscape during training (Foret et al., 2021). Such convergence makes the model vulnerable to slight deviations in the test data, leading to a performance drop. Furthermore, high variability and limited availability of the test data significantly increase the vulnerability of the pre-trained source model. In response, we propose DPO to secure adaptation generalizability and robustness through a worst-case Dual-Perturbation Optimization in both model weight and input spaces. Specifically, at the model level, we apply a perturbation in the weight space (Foret et al., 2021) to the model’s parameters to maximize loss within a predefined range, thereby optimizing the model toward noise-tolerant flat minima. However, due to the notable discrepancies between the training and testing scenes, merely weight perturbation is insufficient to fully address the extensive variability and complexity encountered in the 3D testing scenes. To overcome this, we augment our approach by incorporating an adversarial perturbation on the BEV feature of the test sample via element-wise addition. Once the model is adapted to maintain stability despite perturbed inputs, it becomes more resilient to noisy data, thereby enhancing its robustness. The generalization and robustness of the adaptation model heavily rely on accurate supervision—that is, adapting the detection model based on reliable pseudo-labeled 3D boxes. The supervision signals offered in previous works are either too weak for 3D detection tasks (Wang et al., 2021; Niu et al., 2023) or excessively dependent on pre-trained source models (VS et al., 2023), which might be compromised by domain shifts. To this end, we introduce a reliable Hungarian matcher to ensure trustworthy pseudo-labels by filtering out 3D boxes that exhibit high matching costs before and after perturbations. The underlying assumption is that, given arbitrary perturbations, the prediction is more trustworthy if the model can still produce consistent box predictions. A consistently low Hungarian cost for pseudo-labels across recent test batches indicates the model has been sufficiently robust to shifts/noise in the test domain. Hence, to preserve generalization and minimize unnecessary computational expenses, we propose an Early Hungarian Cutoff strategy based on the Hungarian costs. We apply a moving average of the cost values from the current and all previous batches to determine when to cease the adaptation. Our approach exhibits state-of-the-art results surpassing previous TTA methods. We summarize our key contributions as follows:
  • We introduce TTA in LiDAR-based 3D object detection (TTA-3OD). To the best of our knowledge, this is the first work to adapt the 3D object detector during test time. To tackle the challenge in TTA-3OD, we prioritize the importance of model generalizability and reliable supervision.
  • We propose a dual-perturbation optimization (DPO) mechanism, which maximizes the model perturbation and introduces input perturbation. This strategy is key to maintaining the model’s generalizability and robustness during updates.
  • We leverage a Hungarian matching algorithm to facilitate the selection of noise-insensitive pseudo-labels, to bolster adaptation performance through self-training. This further serves as a criterion for appropriately timing the cessation of model updates.
  • By conducting thorough evaluations of DPO across various scenarios, including cross-domain, corruption-based, and notably complex composite domain shifts, our approach showcases outstanding performance in LiDAR-based 3D object detection tasks, specifically on Waymo $\rightarrow$ KITTI, outperforming the most competitive baseline by 57.72% in $\text{AP}_{\text{3D}}$, and achieve 91% of the fully supervised upper bound.
2. Related Work
2.1. Domain Adaptive 3D Object Detection
Adaptation for 3D Object Detection focuses on transferring knowledge from 3D detectors trained on labeled source point clouds to unlabeled target domains, effectively reducing the domain discrepancies across diverse 3D environments such as variations in object statistics (Wang et al., 2020; Tsai et al., 2023b), weather conditions (Xu et al., 2021; Hahner et al., 2022), sensor differences (Rist et al., 2019; Gu et al., 2021; Wei et al., 2022), sensor failures (Kong et al., 2023), and the synthetic-to-real gap (Saleh et al., 2019; DeBortoli et al., 2021; Lehner et al., 2022). Strategies to overcome these challenges include adversarial feature alignment (Zhang et al., 2021), 3D pseudo-labels (Yang et al., 2021, 2022; Chen et al., 2023; Saltori et al., 2020; Chen et al., 2024; You et al., 2022; Li et al., 2023b; Peng et al., 2023; Huang et al., 2024; Tsai et al., 2023a), the mean-teacher model (Luo et al., 2021; Hegde et al., 2023) for prediction consistency, and contrastive learning (Zeng et al., 2021). Nonetheless, these cross-domain adaptation methods typically necessitate adaptation over multiple epochs, making them less suited for real-time test scenarios.
2.2. Test-time Adaptation in 2D Vision Tasks
Test-time adaptation (TTA) (Liang et al., 2023; Wang et al., 2023c) is designed to address domain shifts between the training and testing data (Wang et al., 2023c) during inference time. As a representative, Tent (Wang et al., 2021) leverages entropy minimization for BatchNorm adaptation. Subsequent works (Hong et al., 2023; Gong et al., 2022; Song et al., 2023; Nguyen et al., 2023; Niloy et al., 2024) such as EATA (Niu et al., 2022), identifies reliable and nonredundant samples to optimize. DUA (Mirza et al., 2022) introduces adaptive momentum in a new normalization layer whereas RoTTA (Yuan et al., 2023) and DELTA (Zhao et al., 2023) leverage global statistics for batch norm updates. Furthermore, SoTTA (Gong et al., 2024) and SAR (Niu et al., 2023) improve BatchNorm optimization by minimizing the loss sharpness. Alternatively, some approaches optimize the entire network through the mean-teacher framework for stable supervision (Wang et al., 2022a; Tomar et al., 2023b), generate reliable pseudo-labels for self-training (Goyal et al., 2022; Zeng et al., 2024), employ feature clustering (Chen et al., 2022; Jung et al., 2023; Wang et al., 2023d), and utilizing augmentations to enhance model robustness (Zhang et al., 2022). However, these TTA methods are developed for general image classification. Additionally, MemCLR (VS et al., 2023) applies TTA for image-based 2D object detection, using a mean-teacher approach to align instance-level features. Nevertheless, the applicability of these image-based TTA methods to object detection from 3D point clouds remains unexplored.
2.3. Generalization through Flat Minima
The concept of flat minima has been demonstrated to enhance model generalization. A prime example is SAM (Foret et al., 2021), which improves generalization by simultaneously optimizing the original objective (e.g., cross-entropy loss) and the flatness of the loss surface. Besides, ASAM (Kwon et al., 2021) aligns the sharpness with the generalization gap by re-weighting the perturbation according to the normalization operator. To deal with the presence of multiple minima within the perturbation’s reach, GSAM (Zhuang et al., 2022) minimizes the surrogate gap between the perturbed and the original loss to avoid sharp minima with low perturbed loss. Moreover, GAM (Zhang et al., 2023) introduces first-order flatness, which controls the maximum gradient norm in the neighborhood of minima. Current research on flat minima focuses mainly on supervised learning. While in TTA, the effectiveness of these strategies significantly relies on supervision signals and the shift severity of the test data, which suggests that the anticipated advantages of flat minima might not consistently materialize as expected.
3. Method
3.1. Notations and Task Definition
Considering a neural network-based 3D object detector $f_{S}(\cdot;\Theta_{\mathcal{S}})$ parameterized by $\Theta_{\mathcal{S}}$, which is pre-trained on a labeled training point clouds drawn from the source distribution $\mathcal{D}_{\mathcal{S}}$, Test-Time Adaptation for 3D Object Detection (TTA-3OD) aims to adapt $f_{\mathcal{S}}(\cdot;\Theta_{\mathcal{S}})$ to the unlabeled test point clouds $\{X_{t}\}_{t=1}^{T}\sim\mathcal{D}_{\mathcal{T}}$ during test time in a single pass. $\mathcal{D}_{\mathcal{S}}\neq\mathcal{D}_{\mathcal{T}}$ as the test point clouds are shifted due to varied real-world conditions. Here, $X_{t}$ represents the $t$-th batch of test point clouds, with $f_{t}(\cdot;\Theta_{t})$ indicating the 3D detection model adapted for the $t$-th batch.
Challenges in TTA-3OD. The primary challenges of TTA-3OD lie in two aspects: (1) adapting the 3D detection model to unfamiliar test scenes often generates large and noisy gradients, leading to an unstable adaptation process. This instability hampers the model’s ability to generalize effectively to the target domain; (2) uncontrollable variations in the testing scenes, such as environmental changes or sensor inaccuracies, can significantly compromise the quality and integrity of 3D scenes. Consequently, models trained on clean datasets struggle to maintain effectiveness and robustness when facing such distorted data, drastically diminishing their adaptation performance.
To address the above two challenges, our method fundamentally enhances 1) the model’s generalization and stability when adapting to new domains and 2) its robustness against noisy/corrupted data, by optimizing the sharpness of the loss landscape during model adaptation with the proposed dual-perturbation applied to both the model’s weights and input data.
3.2. Minimizing Sharpness in the Weight Space
The sharpness of the training loss, is the rate of change in the surrounding region of the loss landscape. It has been identified to be empirically correlated with the generalization error (Ghadimi and Lan, 2013; Jiang et al., 2020; Keskar et al., 2017). Motivated by this, recent works propose to reduce the loss sharpness during the training phase, aiming to improve the generalization capabilities of the model. One notable example is Sharpness-Aware Minimization (SAM), which enhances model training by integrating and optimizing the worst-case perturbations in model weights. The fundamental principle of SAM is that by minimizing the loss with respect to maximally perturbed weights within a vicinity, the entire vicinity (i.e., all losses within it) is minimized. This directs the optimization trajectory toward a flat minima in the loss landscape. A flat minima is indicative of superior generalization capabilities, as the loss over it is less sensitive to large perturbations and/or noise in the model weights, unlike sharp minima.
In the context of TTA-3OD, the loss sharpness (Andriushchenko and Flammarion, 2022) during the adaptation can be formally defined as follows:
Definition 3.0 (Loss Sharpness).
The sharpness of the loss $\mathcal{L}_{\text{det}}(X_{t};\Theta_{t})$ is of 3D detection model $f_{t}(\cdot;\Theta_{t})$ to test the $t$-th batch of target point cloud $X_{t}$, denoted as $s(\Theta_{t},X_{t})$, is given by
(1) $s(\Theta_{t},X_{t})\triangleq\max_{\|\epsilon\|_{2}\leq\rho}\mathcal{L}_{\text{det}}(X_{t};\Theta_{t}+\epsilon)-\mathcal{L}_{\text{det}}(X_{t};\Theta_{t}).$
Here, $\epsilon$ is a perturbation vector in the weight space such that its Euclidean norm is bounded by $\rho$.
Previous literature (Foret et al., 2021; Zhang et al., 2023; Kwon et al., 2021; Du et al., 2022; Zhuang et al., 2022; Liu et al., 2022) calculates the sharpness by the loss between model predictions $f_{t}(X_{t})$ and its ground truth labels $Y_{t}$. While no supervision is available during test time, a soft loss (Gong et al., 2024; Niu et al., 2023; Wang et al., 2021) is commonly employed with selective supervision. Next, the optimization of the detection loss and its sharpness is defined as:
(2) $\begin{split}\min_{\Theta_{t}}\max_{\|\epsilon_{w}\|_{2}\leq\rho}\mathcal{L}_{\text{det}}(X_{t};\hat{Y}_{t};\Theta_{t}+\epsilon_{w}).\end{split}$
The inner optimization aims to find a perturbation $\epsilon_{w}$ on model weights $\Theta_{t}$ within a Euclidean ball of radius $\rho$ to maximize the detection loss $\mathcal{L}_{\text{det}}$, which is calculated based on the generated pseudo-labels $\hat{Y}_{t}$. To obtain the worst-case $\epsilon_{w}$, we draw inspiration from (Foret et al., 2021) to approximate the inner optimization by the first-order Taylor expansion:
(3) $\begin{split}\epsilon_{w}^{*}(\Theta_{t})&\triangleq\underset{\|\epsilon_{w}\|_{2}\leq\rho}{\operatorname*{arg\,max}}\mathcal{L}_{\text{det}}(X_{t};\hat{Y}_{t};\Theta_{t}+\epsilon_{w})\\ &\approx\underset{\|\epsilon_{w}\|_{2}\leq\rho}{\operatorname*{arg\,max}}\ \mathcal{L}_{\text{det}}(X_{t};\hat{Y}_{t};\Theta_{t})+\epsilon_{w}^{\top}\nabla_{\Theta_{t}}\mathcal{L}_{\text{det}}(X_{t};\hat{Y}_{t};\Theta_{t})\\ &=\underset{\|\epsilon_{w}\|_{2}\leq\rho}{\operatorname*{arg\,max}}\ \epsilon_{w}^{\top}\nabla_{\Theta_{t}}\mathcal{L}_{\text{det}}(X_{t};\hat{Y}_{t};\Theta_{t}).\end{split}$
Then $\hat{\epsilon}_{w}(\Theta_{t})$, which satisfies this approximation, is derived by resolving a dual norm problem:
(4) $\begin{split}\hat{\epsilon}_{w}(\Theta_{t})=\rho&\times\text{sign}(\nabla_{\Theta_{t}}\mathcal{L}_{\text{det}}(X_{t};\hat{Y}_{t};\Theta_{t}))\\ &\times\frac{|\nabla_{\Theta_{t}}\mathcal{L}_{\text{det}}(X_{t};\hat{Y}_{t};\Theta_{t})|}{\|\nabla_{\Theta_{t}}\mathcal{L}_{\text{det}}(X_{t};\hat{Y}_{t};\Theta_{t})\|_{2}}.\end{split}$
To expedite computation, we omit the second-order term.
Figure 1. (1) Loss contour for weight perturbation $\hat{\epsilon}_{w}$ (left); (2) The loss profile view for input perturbation $\hat{\epsilon}_{z}$ (right). Our goal is to optimize the loss towards flat minima while ensuring the model’s resilience to data perturbations. Darker colors indicate lower loss values.
While this improves model generalization by targeting non-sharp minima within the loss landscape, the optimized perturbations to weights do not directly deal with variations and/or noise of the input test scenes. When facing the test data, the detection performance is substantially degraded due to the severe data-level corruptions in the test point cloud. For example, when suffering heavy snow and shifted object scales, simultaneously, the performance drops from 73.45% to 3.84% in $\text{AP}_{\text{3D}}$. The empirical evidence suggests that augmenting the baseline model with SAM (Niu et al., 2023) results in a marginal improvement of only 0.9% in $\text{AP}_{\text{3D}}$, indicating its ineffectiveness in bridging the domain gap in the test 3D scenes.
3.3. Minimizing Sharpness in the Input Space
To surmount the above challenge, we strengthen the model’s resilience against shifted input point clouds by optimizing the model with perturbed input. Rather than randomly mimicking test perturbations, our approach focuses on learning an adversarial perturbation that represents the worst-case corruption to the input data. The underlying rationale is that optimizing the detection model with maximal perturbed data within a given vicinity induces robustness to any perturbations encountered within that vicinity. As shown in Figure 1, we simultaneously guide the detection model toward the flat minima in both weight and input space, such that the model can stably generalize to the test data with any potential noises.
To introduce perturbations into the input batch, we incorporate an adversarial perturbation mask $\epsilon_{z}$ into the bird’s eye view (BEV) feature map $Z_{t}$ through element-wise addition to each grid of the BEV map. This is because the 3D detector primarily localizes object proposals from the BEV map, which contains rich spatial information about 3D instances. Thus, even minimal perturbations to the feature map can cause significant spatial shifts in the instances, leading to misalignment in the final predicted 3D bounding boxes. To seek the worst-case perturbation $\epsilon_{z}$ within the input space that maximizes detection loss, we formulate the optimization problem as follows:
(5) $\begin{split}\epsilon_{z}^{*}(Z_{t})&\triangleq\underset{\|\epsilon_{z}\|_{2}\leq\rho}{\operatorname*{arg\,max}}\ \mathcal{L}_{\text{det}}(Z_{t}+\epsilon_{z};\hat{Y}_{t};\Theta_{t}).\end{split}$
Similar to approximating $\hat{\epsilon}_{w}(\Theta_{t})$, we derive the approximated $\hat{\epsilon}_{z}(Z_{t})$ within the input space. This resulting perturbation mask $\hat{\epsilon}_{z}$ shares the same dimension as the latent feature map $Z_{t}$ and is applied to $Z_{t}$ via element-wise addition, yielding the perturbed feature map $Z_{t}+\hat{\epsilon}_{z}$.
The final objective is to train the detection model with the optimal dual-perturbation in both model ($\hat{\epsilon}_{w}$) and input space ($\hat{\epsilon}_{z}$). To this end, we approximate the gradient by substituting $\hat{\epsilon}_{w}$ and $Z_{t}+\hat{\epsilon}_{z}$ into Eqn. (2), then performing differentiation to calculate the gradient $g$:
(6) $\begin{split}g=\nabla_{\Theta_{t}}\mathcal{L}_{\text{det}}(Z_{t}+\hat{\epsilon}_{z};\hat{Y}_{t};\Theta_{t})|_{\Theta_{t}+\hat{\epsilon}_{w}}.\end{split}$
Finally, the detection loss and its sharpness, calculated with the perturbed test batch, can be jointly minimized by:
(7) $\begin{split}\min_{\Theta_{t}}\max_{\begin{subarray}{c}\|\epsilon_{w}\|_{2}\leq\rho\\ \|\epsilon_{z}\|_{2}\leq\rho\end{subarray}}\mathcal{L}_{\text{det}}(Z_{t}+\epsilon_{z};\hat{Y}_{t};\Theta_{t}+\epsilon_{w}),\end{split}$
where the inner optimization is solved through approximation (i.e., Eqn.  (3)–(5)) and the outer optimization goal is achieved by stochastic gradient descent (SGD) with the gradient $g$ calculated in Eqn. (6). The step-by-step workflow of the proposed DPO is introduced in Algorithm 1.
3.4. Reliable Hungarian Matcher
However, the pursuit of flat minima in both the input and weight spaces depends on the gradients guided by high-quality supervision. Previous SAM-based TTA methods selectively adapt high-confidence samples (Gong et al., 2024; Niu et al., 2023; Yu et al., 2023), as they assume that confidence reflects prediction reliability. Nevertheless, acquiring effective supervision in the TTA-3OD task is challenging due to the low-quality pseudo-labeled boxes, $\hat{Y}_{t}=\{\hat{b}_{1},\cdots,\hat{b}_{N_{t}}\}$, used for calculating the detection loss $\mathcal{L}_{\text{det}}$, where $N_{t}$ represents the number of predicted boxes in the current batch $t$. This issue arises from the difficulties the source-trained model $f_{S}(\cdot;\Theta_{\mathcal{S}})$ faces in accurately predicting 3D boxes around objects in the test point clouds, which subjects to significant shifts or corruptions.
Figure 2. Illustration of the proposed Hungarian matcher for obtaining reliable supervision. We employ the Hungarian algorithm to compute the cost for each pseudo-labeled 3D box (i.e., predictions before perturbation) when paired with its optimally matched counterpart in predictions after perturbation. The reliability of the 3D boxes is categorized into three tiers—high, medium, and low—based on the computed matching cost. During TTA, only 3D boxes of high reliability (e.g., ID 1, 2) are used for updating model weights, and those of low reliability (e.g., ID 4) are treated as background.
To obtain reliable pseudo-labeled boxes that are robust to the test data noise, we aim to select those 3D boxes unaffected by optimized perturbations (Sect. 3.3). The rationale is that consistency in box predictions between clean inputs and perturbed input features $Z_{t}+\hat{\epsilon}_{z}$ from the model before and after perturbation demonstrates resilience to noises $\hat{\epsilon}_{z}$. The box prediction from the $t$-th perturbed input batch is defined as:
(8) $\begin{split}\tilde{Y}_{t}=\{\tilde{b}_{1},\cdots,\tilde{b}_{M_{t}}\}=f_{t}(Z_{t}+\hat{\epsilon}_{z};\Theta_{t}+\hat{\epsilon}_{w}),\end{split}$
where $M_{t}$ is the number of predicted boxes at batch $t$ after perturbation. To measure the consistency between $\hat{Y}_{t}$ and $\tilde{Y}_{t}$, we adopt Hungarian matching (Carion et al., 2020; Stewart et al., 2016), an effective bipartite matching technique that guarantees optimal one-to-one alignment between two sets of box predictions. Specifically, We ensure both sets are of equal size by augmenting the smaller set (assuming $M_{t}<N_{t}$) with $\emptyset$ until it matches $N_{t}$ in size. To achieve optimal bipartite matching between the equal-sized sets, the Hungarian algorithm is applied to find a permutation of $N_{t}$ elements $p\in\mathbf{P}_{N_{t}}$ that minimizes the matching cost:
(9) $\tilde{p}=\operatorname*{arg\,min}_{p\in\mathbf{P}_{N_{t}}}\sum_{n}^{N_{t}}\mathcal{C}_{\text{box}}(\hat{b}_{n};\tilde{b}_{p(n)}).$
The cost $\mathcal{C}_{\text{box}}(\cdot;\cdot)$ integrates intersection-over-union (IoU) and L1 distance to account for the central coordinates, dimensions, and orientations of a pair of boxes: $\hat{b}_{n}$ and its corresponding matched box $\tilde{b}_{p(n)}$ indexed by $p(n)$.
Utilizing the derived optimal assignment $\tilde{p}$, each pseudo-labeled box $\hat{b}_{n}$ is associated with its corresponding minimal-cost match, denoted as $\mathcal{C}_{\text{box}}(\hat{b}_{n};\tilde{b}_{\tilde{p}(n)})$. Note that when $\hat{b}_{n}$ pairs with the empty set $\emptyset$, we assign the cost to be infinite, indicating that the corresponding box is too noise-sensitive to be accurately localized within the perturbed point clouds.
We categorize the reliability of each pseudo-label $\hat{b}_{n}$ into three distinct levels by the thresholds $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$:
[…]
To guide the model towards flat minima with trustworthy supervision, we selectively adapt the model with high-quality bounding boxes and treat those of low quality as background, as shown in Figure 2. To dynamically set the thresholds $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, we record the minimum costs of pseudo-labeled boxes from previous batches in a sorted array $A_{\text{costs}}$, then determine $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ as the upper and lower $\alpha$ quantiles:
(11) $\begin{split}\mathcal{C}_{1}=A_{\text{costs}}[\lceil\alpha\times n\rceil],\mathcal{C}_{2}=A_{\text{costs}}[\lceil(1-\alpha)\times n\rceil],\\ \text{where}\ A_{\text{costs}}=\text{sort}(\{{\mathcal{C}_{\text{box}}(\hat{b}_{n};\tilde{b}_{\tilde{p}(n)})}\}),\\ \hat{b}_{n}\in\{\hat{Y}_{1}\}\cup\cdots\{\hat{Y}_{t}\},\tilde{b}_{n}\in\{\tilde{Y}_{1}\}\cup\cdots\{\tilde{Y}_{t}\}.\end{split}$
[…]
3.5. Early Hungarian Cutoff
While the Hungarian matcher significantly enhances the quality of pseudo-labels, the correctness of the selected pseudo-labels cannot be entirely guaranteed. Even a small number of incorrect pseudo-labels once learned and accumulated, can lead to substantial performance degradation. Furthermore, updating the 3D detector demands significant computational resources and time. Identifying an optimal stopping point for the adaptation process is thus crucial.
In this regard, we suggest using the Hungarian cost as a criterion to halt the adaptation process. The rationale is that a lower Hungarian cost for a given batch indicates the pseudo-labels are more accurate, thereby making the update process more reliable. Additionally, a consistently low Hungarian cost of pseudo-labels is crucial. Therefore, we introduce the use of a moving average to balance the current and all previous costs:
(12) $\begin{split}C_{\text{ema}}^{t}=\gamma C_{\text{box}}^{t}+(1-\gamma)\sum C_{\text{box}}^{t-1},\\ \text{where }C_{\text{box}}^{t}=\frac{1}{N_{t}}\sum_{n}^{N_{t}}\mathcal{C}_{\text{box}}(\hat{b}_{n};\tilde{b}_{\tilde{p}(n)}),\hat{b}_{n}&\in\hat{Y}_{t},\tilde{b}_{n}\in\tilde{Y}_{t},\end{split}$
is the average Hungarian cost of all boxes in the current batch $t$. $C_{\text{ema}}^{t}$ represents the moving average of the Hungarian cost, and $\gamma$ denotes the decay rate. A threshold $C_{\text{stop}}$ is further set for the moving average cost. When it falls below the threshold, the adaptation process is halted, the model thus transitions to the inference mode for all subsequent batches.
Algorithm 1 DPO for TTA-3OD
$f_{S}(\cdot;\Theta_{\mathcal{S}})$: source pre-trained model, $\{X_{t}\}_{t=1}^{T}\sim\mathcal{D}_{\mathcal{T}}$: target point clouds to test, $\eta$: step size, $C_{\text{stop}}$: early-stop threshold
$f_{t}(\cdot;\Theta_{t})$: model adapted to the target point clouds.
Initiate the weights $\Theta_{1}=\Theta_{\mathcal{S}}$
for $t=1,\cdots,T$ do
     Generate predictions $\hat{Y}_{t}\leftarrow f_{t}(X_{t};\Theta_{t})$ as pseudo-label
     Compute perturbations $\hat{\epsilon}_{z}$, $\hat{\epsilon}_{w}$ via Eqn. (3)–(5)
     Generate prediction $\tilde{Y}_{t}$ with perturbations via Eqn. (8)
     Refine $\hat{Y}_{t}$ by reliable Hungarian matcher with $\tilde{Y}_{t}$ via  (9)–(11)
     Compute gradient approximation $g$ via Eqn. (6)
     Update weights: $\Theta_{t+1}=\Theta_{t}-\eta g$
     /** check early stopping **/
     Compute the Hungarian matching cost $C_{\text{ema}}^{t}$ via Eqn. (12)
     if $C_{\text{ema}}^{t}\leq C_{\text{stop}}$ then break
         Infer the remaining batches with $f_{t}(X_{t};\Theta_{t})$
     end if
end for
Table 1. Results of test-time adaption to 3D scenes under cross-dataset shift. We report $\text{AP}_{\text{BEV}}$ / $\text{AP}_{\text{3D}}$ at moderate difficulty. Oracle means fully supervised training on the target dataset. The best adaptation results are highlighted in bold.
Method Venue TTA Waymo →KITTI nuScenes →KITTI
[…]
No Adapt. - - 67.64 / 27.48 - 51.84 / 17.92 -
SN CVPR’20 $\times$ 78.96 / 59.20 $+$72.33% / $+$69.00% 40.03 / 21.23 $+$37.55% / $+$5.96%
ST3D CVPR’21 $\times$ 82.19 / 61.83 $+$92.97% / $+$74.72% 75.94 / 54.13 $+$76.63% / $+$65.21%
Oracle - - 83.29 / 73.45 - 83.29 / 73.45 -
Tent ICLR’21 65.09 / 30.12 --16.29% / $+$5.74% 46.90 / 18.83 --15.71% / $+$1.64%
CoTTA CVPR’22 67.46 / 35.34 --1.15% / $+$17.10% 68.81 / 47.61 $+$53.96%/ $+$53.47%
SAR ICLR’23 65.81 / 30.39 --11.69% / $+$6.33% 61.34 / 35.74 $+$30.21% / $+$32.09%
MemCLR WACV’23 65.61 / 29.83 --12.97% / $+$5.11% 61.47 / 35.76 $+$30.62% / $+$32.13%
DPO - 75.81 / 55.74 $+$52.20% / $+$61.47% 73.27 / 54.38 $+$68.13%/$+$65.66%
4. Experiments
4.1. Experimental Setup
4.1.1. Datasets and TTA-3OD Tasks.
Our experiments are carried out on three widely used LiDAR-based 3D object detection datasets: KITTI (Geiger et al., 2012), Waymo (Sun et al., 2020), and nuScenes (Caesar et al., 2020). Additionally, the recently released KITTI-C dataset (Kong et al., 2023), which simulates real-world corruptions, is incorporated to pose the TTA-3OD challenge. Following prior works (Yang et al., 2021, 2022; Chen et al., 2023), we address cross-dataset test-time adaptation tasks (e.g., Waymo $\rightarrow$ KITTI and nuScenes $\rightarrow$ KITTI), involving adaptation across (i) object shifts (e.g., scale and point density variations), and (ii) environmental shifts (e.g., changes in deployment locations and LiDAR configurations). Furthermore, we evaluate adaptation performance against real-world corruptions (e.g., KITTI $\rightarrow$ KITTI-C), including conditions such as fog, wet conditions (Wet.), snow, motion blur (Moti.), missing beams (Beam.), crosstalk (Cross.T), incomplete echoes (Inc.), and cross-sensor interference (Cross.S). Experiments also extend to the challenging scenarios of composite domain shifts (e.g., Waymo $\rightarrow$ KITTI-C), where inconsistencies across datasets and corruptions coexist within the test 3D scenes.
4.1.2. Implementation Details.
We leverage the OpenPCDet framework (Team, 2020). Experiments are conducted on a single NVIDIA RTX A6000 GPU with 48 GB of memory. We opt for a batch size of $8$ and fix the hyperparameters $\rho=1e-4,\alpha=0.08,\gamma=0.5,\eta=10^{-3}$. For evaluation purposes, we adhere to the official metrics of the KITTI benchmark, reporting the average precision for the car class in both 3D (i.e., $\text{AP}_{\text{3D}}$) and bird’s eye view (i.e., $\text{AP}_{\text{BEV}}$) perspectives, calculated over 40 recall positions and applying a $0.7$ IoU threshold. The closed gap (Yang et al., 2021) is calculated as: $\frac{\mathrm{AP}_{\text{method }}-\mathrm{AP}_{\text{No Adapt. }}}{\mathrm{AP}_{\text{Oracle }}-\mathrm{AP}_{\text{No Adapt. }}}\times 100\%$.
4.1.3. Baseline Methods.
We integrate a voxel-based backbone (i.e., SECOND) into our proposed method and evaluate it against a comprehensive array of baseline approaches:
  • No Adapt.: directly inferring the test data with a model pre-trained on the source domain, without any adaptation.
  • SN (Wang et al., 2020): weakly supervised domain adaptive 3D detection that adjusts source object sizes using target domain statistics.
  • ST3D (Yang et al., 2021): an unsupervised domain adaptation method for 3D detection, utilizing multi-epoch pseudo-labeling for self-training.
  • Tent (Wang et al., 2021): a fully TTA method that optimizes BatchNorm layers by minimizing the entropy of predictions.
  • CoTTA (Wang et al., 2022a): a TTA strategy that leverages mean-teacher framework to provide supervisory signals through augmentations and employs random neuron restoration to retain source knowledge.
  • SAR (Niu et al., 2023): an advancement beyond Tent, employing sharpness-aware minimization for selected supervision.
  • MemCLR (VS et al., 2023): TTA for image-based object detection using mean-teacher to align the instance-level features by a memory module.
  • Oracle: a fully supervised model trained on the test scenes.
4.2. Experimental Results
4.2.1. Cross-dataset Shifts.
We conducted extensive experiments on two cross-dataset TTA-3OD tasks, evaluating $\text{AP}_{\text{BEV}}$, $\text{AP}_{\text{3D}}$, and closed gap, as presented in Table 1. Compared to direct inference (i.e. No Adapt.), our experiments revealed that existing TTA baselines might negatively impact adaptation in 3D object detection especially on $\text{AP}_{\text{BEV}}$ for the Waymo $\rightarrow$ KITTI task, indicating the importance of tailoring a TTA method specifically for 3D detection tasks. Additionally, compared to the most competitive baseline, CoTTA, DPO achieves significant improvements in $\text{AP}_{\text{3D}}$, with increases of 57.7% and 14.2% for the Waymo $\rightarrow$ KITTI and nuScenes $\rightarrow$ KITTI tasks, respectively. Similarly, DPO significantly outperforms CoTTA in $\text{AP}_{\text{BEV}}$, demonstrating a considerable margin. Besides, DPO effectively reduces the closed gap, demonstrating a closure of about 61.47% and 65.66% for the Waymo $\rightarrow$ KITTI and nuScenes $\rightarrow$ KITTI tasks, correspondingly, in $\text{AP}_{\text{3D}}$. Moreover, it achieves up to 91% and 87.5% of the fully supervised Oracle’s performance in $\text{AP}_{\text{BEV}}$ for the respective tasks. Overall, our proposed DPO not only surpasses all TTA baselines but also delivers performances competitive with those of Unsupervised Domain Adaptation (UDA) and fully supervised learning, highlighting its effectiveness in bridging domain gaps in 3D object detection.
Table 2. Results of KITTI $\rightarrow$ KITTI-C on heavy corruptions.
No Adapt. Tent CoTTA SAR MemCLR DPO
Fog 68.23 68.73 68.49 68.14 68.23 68.72
Wet. 76.25 76.36 76.43 76.23 76.25 76.89
Snow 59.07 59.50 59.45 58.78 58.74 60.80
Moti. 38.21 38.15 38.62 38.12 37.57 38.71
Beam. 53.93 53.85 53.98 53.75 53.49 54.06
CrossT. 75.49 74.67 72.22 74.51 74.25 75.52
Inc. 25.68 26.44 27.35 26.42 27.47 27.16
CrossS. 41.08 41.17 40.80 40.63 40.90 42.09
Mean 54.74 54.86 54.67 54.57 54.61 55.49
4.2.2. Corruption Shifts.
We evaluated DPO’s efficacy against corruption-induced shifts on KITTI $\rightarrow$ KITTI-C with heavy severity of eight real-world corruption by $\text{AP}_{\text{3D}}$ in hard difficulty scenarios. As indicated in Table 2, DPO outperforms all TTA baselines in terms of Mean $\text{AP}_{\text{3D}}$, exceeding the performance of the closest competitive baseline, Tent, by 1.2%. DPO consistently achieves top performance across most corruption types, demonstrating the enhanced robustness of DPO and its effectiveness in adapting 3D models to a wide array of corrupted environments.
Figure 3. Results ($\text{AP}_{\text{3D}}$) of adapting across composite shifts (Waymo $\rightarrow$ KITTI-C) at the heavy corruption level. Lighter shades indicate higher performance.
4.2.3. Composite Domain Shifts.
To address the most challenging shift in 3D scenes, which merges both cross-dataset discrepancies and corruptions, we conducted experiments to adapt 3D detectors from Waymo to KITTI-C (heavy corruption). The outcomes are represented in Figure 3. Notably, the shades in the last column (DPO) are significantly lighter than those in all other columns (TTA baselines) at various difficulty levels (moderate and hard), indicating DPO’s superior performance. In particular, the performance without any adaptation (column 1) significantly declines, illustrating the compounded challenges of composite shifts. For example, only 8.38% $\text{AP}_{\text{3D}}$ is recorded for Motion Blur and 3.84% $\text{AP}_{\text{3D}}$ for Incomplete Echoes at the moderate level. Conversely, against the most competitive baseline (column 3), our approach notably enhances adaptation performance for these challenging corruptions by 75.13% and 61.36%, respectively. Direct inference for Incomplete Echoes at hard difficulty yields only a 3.51% in $\text{AP}_{\text{3D}}$, whereas our method markedly increases this by more than 231.62%, achieving a 53.97% improvement over the highest baseline. In summary, existing TTA methods fall short in navigating significant domain shifts (i.e., composite domain shifts) in 3D scenes, while DPO could effectively tackle these challenges.
4.3. Parameter Sensitivity
Figure 4. Sensitivity to radius $\rho$ in SAM (left), and the pseudo-label threshold $\alpha$ (right) on nuScenes $\rightarrow$ KITTI.
4.3.1. Sharpness Radius $\rho$.
To understand the impact of varying the sharpness radius $\rho$ on $\text{AP}_{\text{3D}}$ and $\text{AP}_{\text{BEV}}$, we conduct an analysis at the moderate difficulty level for the nuScenes $\rightarrow$ KITTI task, keeping all other hyperparameters fixed. We explored a range of $\rho$ values from $10^{-4}$ to $10^{-1}$. The left part of Figure 4 illustrates that variations in $\rho$ exhibit a minimal influence on $\text{AP}_{\text{BEV}}$, contrasting with $\text{AP}_{\text{3D}}$, which demonstrates significant performance variability when the perturbation radius is adjusted to $0.1$. This discrepancy can be attributed to two primary factors. Firstly, an increase in perturbation radius adversely affects adaptation performance. Secondly, a larger perturbation radius results in the selection of a reduced number of pseudo-labeled 3D boxes for self-training due to the increased divergence in model predictions. However, when employing a perturbation radius within a smaller range (e.g., $10^{-4}$-$10^{-2}$), the stability of $\text{AP}_{\text{3D}}$ is notably enhanced.
4.3.2. Pseudo-label Threshold $\alpha$
The pseudo-label threshold $\alpha$ shows a consistent pattern for $\text{AP}_{\text{BEV}}$, remaining stable across different values. However, a low threshold (i.e., 0.04) causes a drop in $\text{AP}_{\text{3D}}$ as too few pseudo-labeled 3D boxes are selected to update model weights. This emphasizes the need for an appropriate proportion of pseudo-labels for adaptation. For $\alpha$ values between 0.08 and 0.20, $\text{AP}_{\text{BEV}}$ and $\text{AP}_{\text{3D}}$ remain stable, with maximum fluctuations of 0.83 and 0.07, respectively. This stability highlights the robustness of the selected threshold.
4.4. Ablation Study
Table 3. Ablation study on the nuScenes $\rightarrow$ KITTI task. $\text{AP}_{\text{BEV}}$ (left) and $\text{AP}_{\text{3D}}$ (right) (%) are reported for three levels of difficulty. The best results are highlighted in bold.
Pert. $\Theta_{t}$ Pert. $Z_{t}$ Matcher Easy Moderate Hard
- - - $76.51/58.78$ $62.68/43.64$ $59.93/39.87$
$\surd$ - - $82.14/56.36$ $70.86/47.18$ $68.91/44.62$
$\surd$ $\surd$ - $80.42/60.50$ $72.42/49.28$ $70.77/46.20$
$\surd$ - $\surd$ $81.08/62.88$ $73.09/49.60$ $71.86/47.19$
$\surd$ $\surd$ $\surd$ 83.11/66.19 73.27/54.38 72.21/52.66
4.4.1. Impact of Components.
To understand how individual components of DPO affect overall performance, we conduct an ablation study by incrementally adding each component to adaptation and evaluating performance on the nuScenes $\rightarrow$ KITTI task. Table 3 shows the impact of these components on the KITTI dataset at three difficulty levels, measured by AP score. Here, Pert. $\Theta_{t}$ represents weight space perturbation, Pert. $Z_{t}$ denotes input perturbation, and Matcher refers to the Hungarian Matching mechanism for pseudo-label selection. Compared to the self-training baseline (row 1), adding weight space perturbation (row 2) significantly improves $\text{AP}_{\text{BEV}}$ but reduces $\text{AP}_{\text{3D}}$ (58.78 $\rightarrow$ 56.36 at easy difficulty), indicating limitations of SAM for the TTA-3OD task. Incorporating input perturbation (row 3) and using dual perturbation improves performance, increasing $\text{AP}_{\text{BEV}}$ and $\text{AP}_{\text{3D}}$ across all difficulty levels. The Hungarian matcher enhances pseudo-label selection with weight perturbation alone, as shown by the performance gains over weight perturbation alone (rows 2, 4). Using all proposed DPO components yields the highest performance for both $\text{AP}_{\text{BEV}}$ and $\text{AP}_{\text{3D}}$ across all difficulty levels.
4.4.2. Impact of Early Hungarian Cutoff.
Figure 5. Performance trend and variation in the number of point clouds for weight updates across different early-stopping thresholds $C_{\text{stop}}$.
We examine the effectiveness of early Hungarian cutoff on the nuScenes $\rightarrow$ Waymo task in Figure 5. When the Hungarian cost in Eqn. (12) falls below a specified threshold, e.g., $C_{\text{stop}}=4.8$, the model updates its weights using self-training on the first 48 test point clouds and then infers the remaining point clouds directly, skipping further self-training. The moving-average Hungarian cost converges rapidly during self-training. For instance, using the first 64 test samples reduces the cost from 4.8 to 3.8 and significantly improves performance (9.37 in $\text{AP}_{\text{3D}}$). In contrast, reducing the cost from 3.6 to 3.4 with 144 test samples only marginally improves performance (3.53 in $\text{AP}_{\text{3D}}$) due to error accumulation in pseudo-labels. These results underscore the value of the Hungarian cost-based early stopping mechanism, which leverages a small portion of test batches to enhance performance without excessive computational cost.
4.4.3. Impact of Updating Strategies.
We explore updating only the BatchNorm (BN) vs. the full model for adaptation on Waymo $\rightarrow$ KITTI. As shown in Table 4, updating BN (only 2% of the total parameters) results in a slight decrease of 0.41% in $\text{AP}_{\text{3D}}$ and a slight increase in speed by 0.02s per frame. This demonstrates that our method remains effective even when only a small fraction of the parameters are updated.
Table 4. DPO variations for model updating strategies.
##\## of Params Speed $\text{AP}_{\text{BEV}}$ $\text{AP}_{\text{3D}}$
Full 12,182,565 0.33s / frame 75.81 55.74
BatchNorm 268,288 0.31s / frame 75.12 55.51
4.4.4. Sensitivity to 3D Backbone Detector.
Table 5. DPO results of Waymo $\rightarrow$ KITTI using PV-RCNN.
TTA Method No Adapt. Tent CoTTA SAR Mem-CLR Ours
[…]
4.5. Efficiency Analysis
Figure 6. Efficiency analysis of Waymo $\rightarrow$ KITTI task.
To assess the efficiency of DPO, we conducted a comparative analysis for adaptation speed (i.e., seconds per frame) and GPU memory usage, as illustrated in Figure 6. CoTTA is identified as the most efficient TTA baseline for 3D object detection, demonstrating rapid adaptation capabilities (under 0.15 seconds per frame). Conversely, other baselines, notably SAR, required significantly more adaptation time and yet underperformed, achieving $\text{AP}_{\text{BEV}}$ of less than 66%. Despite a slightly longer processing time, DPO markedly surpassed all TTA baselines, showcasing its superior performance. In terms of GPU memory consumption, CoTTA reported moderate usage, whereas MemCLR exceeded 18,000 MiB but fell short in performance. The proposed DPO, in contrast, not only required less GPU memory than both MemCLR and CoTTA but also achieved dominating adaptation performance, highlighting the efficiency and effectiveness of our method.
5. Conclusion
In this work, we present a novel framework for Test-Time Adaptation in 3D Object Detection (TTA-3OD) aimed at adapting detectors to new unlabeled scenes with a single pass. Our approach incorporates worst-case perturbations at both model and input levels to enhance robustness and generalization, thereby enabling 3D detectors to stably adapt to any test scenes with corruptions. We employ reliable Hungarian matching for trustworthy pseudo-label selection, with an early cutoff to avoid computation burden and error accumulation. Beyond point- and voxel-representation-based 3D detectors used in this paper, our future work will further validate multimodal detectors with different input modalities, such as BEVfusion (Liu et al., 2023) to verify shifts across modalities.
Acknowledgements.
This research is partially supported by the Australian Research Council (DE240100105, DP240101814, DP230101196, DP230101753).
Supplementary Materials
This supplementary material provides additional descriptions of the proposed DPO, including empirical results and implementation details. Visual aids are also included to enhance understanding of the method. Furthermore, the attached code is available for reference.
  • Sect. A: Additional experimental results.
  • Sect. B: More implementation details.
  • Sect. C: Quantitative study on Waymo $\rightarrow$ KITTI-C task.
Table 6. TTA-3OD results (easy/moderate/hard $\text{AP}_{\text{3D}}$) of pedestrian class under the composite domain shift (Waymo $\rightarrow$ KITTI-C) at heavy corruption level.
No Adaptation Tent CoTTA SAR MemCLR DPO
Fog 30.48/26.15/23.61 31.22/26.68/23.99 31.29/26.69/24.05 30.68/25.94/23.70 30.51/26.02/23.77 33.25/27.83/25.23
Wet. 49.10/44.44/41.74 49.13/44.58/41.85 49.14/45.01/42.23 49.18/44.59/41.97 49.09/44.55/41.81 50.31/45.27/42.31
Snow 47.22/42.26/39.19 47.55/42.79/39.44 46.30/41.62/38.11 47.42/42.85/39.54 47.68/42.78/39.45 48.06/43.61/40.11
Moti. 27.18/25.02/23.29 27.47/25.25/23.43 27.28/25.43/23.41 27.34/25.15/23.31 27.44/25.19/23.36 27.48/25.57/23.60
Beam. 32.47/27.89/25.27 34.50/30.55/28.18 32.22/27.41/25.13 34.83/30.74/28.54 34.53/30.30/28.16 34.29/30.42/28.13
CrossT. 47.42/43.08/40.37 47.66/43.37/40.51 47.76/43.29/40.39 48.13/43.65/40.71 47.87/43.58/40.48 50.38/45.43/42.06
Inc. 49.28/44.79/42.21 49.18/44.80/42.11 49.36/45.39/42.77 49.22/44.70/42.24 49.01/44.76/42.11 50.83/46.06/43.02
CrossS. 22.46/18.40/16.08 27.70/22.82/20.30 22.11/17.88/15.98 27.99/23.20/21.36 27.23/23.70/21.63 25.32/20.93/18.99
Mean 38.20/34.00/31.47 39.30/35.11/32.48 38.18/34.09/31.52 39.35/35.10/32.67 39.17/35.11/32.60 39.99/35.64/32.93
No Adaptation Tent CoTTA SAR MemCLR DPO
Fog 21.15/17.91/16.66 23.62/19.21/18.33 22.60/18.74/17.57 23.49/19.02/18.10 23.43/19.01/17.86 23.83/19.61/18.64
Wet. 60.36/49.61/47.20 59.72/48.57/45.96 61.36/49.27/47.04 57.43/46.48/43.96 57.76/46.34/44.79 62.64/50.78/48.50
Snow 48.87/40.37/37.96 52.81/42.25/40.11 52.91/41.55/39.09 52.09/41.89/39.26 52.24/41.71/39.28 58.02/44.09/42.18
Moti. 34.62/29.25/27.33 36.79/29.04/27.31 40.37/31.18/29.34 37.53/29.78/28.12 38.19/29.75/28.19 44.03/34.23/32.05
Beam. 32.48/22.42/21.26 36.08/25.03/23.85 30.89/21.37/20.42 37.16/26.21/24.74 36.34/25.35/24.32 38.65/26.78/25.45
CrossT. 59.56/48.75/46.20 58.72/49.26/46.51 62.14/49.13/46.21 58.66/49.16/46.61 58.68/48.87/46.40 63.57/51.07/48.22
Inc. 59.62/49.03/46.82 59.14/47.87/45.11 59.89/47.62/45.44 58.86/47.93/45.51 58.91/48.28/45.68 62.34/50.18/47.73
CrossS. 18.38/11.40/10.93 24.84/15.04/14.66 20.98/12.77/12.28 26.19/15.29/14.95 25.46/15.06/14.49 24.23/14.28/13.99
Mean 41.88/33.59/31.79 43.96/34.53/32.73 43.89/33.95/32.17 43.93/34.47/32.66 43.88/34.30/32.63 47.16/36.38/34.60
Appendix A Additional Experimental Results
We adhere to a classical LiDAR-based 3D object detection evaluation, focusing on the car class in the main paper. In addition to this, we also explore the effectiveness of the proposed DPO on two other classes: pedestrians and cyclists. We evaluate TTA baselines and DPO across all difficulty levels for the most challenging transfer task, i.e., composite domain shift, in terms of $\text{AP}_{\text{3D}}$. Detailed explanations are provided below.
A.1. Pedestrian Class
We evaluate the effectiveness of DPO for the pedestrian class across all difficulty levels, as shown in Table 6. Notably, our proposed method achieves state-of-the-art performance in terms of mean $\text{AP}_{\text{3D}}$, showcasing its effectiveness. When examining specific corruption types, DPO also demonstrates competitive performance. Specifically, for the crosstalk (CrossT.) corruption, DPO improves the $\text{AP}_{\text{3D}}$ from 40.71% to 42.06%, compared with the strongest baseline SAR, at the hard level. Moreover, our method achieves a 4.1% improvement at the moderate level for the same corruption type. However, there are two exceptions: Beam missing (Beam.) and cross sensor (CrossS.), where a significant number of object points are dropped when generating the corruption, leading to a performance decline for all pseudo-label-based adaptation methods (Kong et al., 2023), including both CoTTA and DPO. Despite these challenges, our method still manages to handle most corruptions effectively, maintaining a leading mean $\text{AP}_{\text{3D}}$.
A.2. Cyclist Class
A similar performance trend is observed in the cyclist class. Our method outperforms the baseline methods at every difficulty level, except for the cross-sensor corruption. Specifically, in terms of mean $\text{AP}_{\text{3D}}$, DPO achieves 7.28%, 5.09%, and 5.71% for the easy, moderate, and hard difficulty levels, respectively. Notably, for the snow corruption, our method leads to the greatest improvement over the baseline, increasing from 52.91% to 58.02% $\text{AP}_{\text{3D}}$ at the easy level. Similarly, a performance increase from 29.34% to 32.05% is achieved at the hard level when facing motion blur (Moti.). For reasons similar to those discussed in Sect. A.1, DPO underperforms for cross-sensor (CrossS.) corruption, potentially due to failure of the pseudo-labeling strategy when encountering cyclists with too few points. However, our method represents the best trade-off solution, as it offers the highest mean $\text{AP}_{\text{3D}}$.
Appendix B More Implementation Details
B.1. Datasets
B.1.1. Waymo
The Waymo open dataset (Sun et al., 2020) is a large 3D detection dataset for autonomous driving. It contains 798 training sequences with 158,361 LiDAR samples and 202 validation sequences with 40,077 LiDAR samples. The point clouds feature 64 lanes of LiDAR, corresponding to 180,000 points every 0.1 seconds. In DPO, we train the source model on the Waymo training set.
B.1.2. nuScenes
The nuScenes dataset (Caesar et al., 2020) consists of 1,000 driving sequences, divided into 700 for training, 150 for validation, and 150 for testing. Each sequence is approximately 20 seconds long, with a LiDAR frequency of 20 FPS. The dataset provides calibrated vehicle pose information for each LiDAR frame while offering box annotations every ten frames (0.5s). nuScenes uses a 32-lane LiDAR, which generates approximately 30,000 points per frame. In total, there are 28,000 annotated frames for training, 6,000 for validation, and 6,000 for testing. We employ its training set for pre-training the source model for all baselines and the proposed DPO.
B.1.3. KITTI
The KITTI Dataset (Geiger et al., 2012) is widely recognized as a crucial resource for 3D object detection in autonomous driving. The training point clouds are divided into a training split of 3,712 samples and a validation split of 3,769 samples. The dataset categorizes detection difficulty into three levels, defined by criteria of visibility, occlusion, and truncation. The category ‘Easy’ denotes scenarios with no occlusion and a truncation limit of 15%. ‘Moderate’ applies to conditions with partial occlusion and truncation not exceeding 30%. ‘Hard’ encompasses situations with severe occlusion and a truncation threshold of 50%. For evaluating the predicted boxes in 3D object detection, KITTI requires a minimum 3D bounding box overlap of 70% for cars and 50% for pedestrians and cyclists. In this study, where KITTI serves as the target domain, we evaluate all models using the validation split.
B.1.4. KITTI-C
The robustness of 3D perception systems against natural corruptions, which arise due to environmental and sensor-related anomalies, is crucial for safety-critical applications. While existing large-scale 3D perception datasets are often meticulously curated to exclude such anomalies, this does not accurately represent the operational reliability of perception models. KITTI-C (Kong et al., 2023) is the first comprehensive benchmark designed to assess the robustness of 3D detectors in scenarios involving out-of-distribution natural corruptions encountered in real-world environments. It specifically investigates three major sources of corruption likely to impact real-world deployments: 1) severe weather conditions such as fog, rain (Wet.), and snow, which affect laser pulse dynamics through back-scattering, attenuation, and reflection; 2) external disturbances including bumpy surfaces, dust, and insects, which can cause motion blur (Moti.) and missing LiDAR beams (Beam.); and 3) internal sensor failures like incomplete echo (Inc.) or misidentification of dark-colored objects and sensor crosstalk (Cross.T), which may compromise 3D perception accuracy. Additionally, understanding cross-sensor discrepancies is essential to mitigate risks associated with sudden failures due to changes in sensor configurations (Cross.S).
B.2. Additional Implementation Details
Tent and SAR (Wang et al., 2021; Niu et al., 2023) utilize entropy minimization to optimize the batchnorm layers during test time. Therefore, we calculate the entropy loss by summing the classification logits for all proposals at the first detection stage. CoTTA (Wang et al., 2022a) follows a mean-teacher framework. Although a broad range of data augmentations is typically required to generalize the model to various corruptions, our empirical evidence from test-time adaptation for 3D object detection (TTA-3OD) suggests that most augmentations do not improve—and may even impair—performance. The sole exception is random world scaling. As a result, we adopt random world scaling as our primary strategy, in accordance with (Luo et al., 2021), applying strong scaling (0.9 to 1.1) and weak scaling (0.95 to 1.05), respectively. Regarding pseudo-labeling, we directly apply strategies tailored for 3D object detection from (Yang et al., 2021, 2022) to enhance self-training in CoTTA. Similar to CoTTA, we adopt the same augmentation strategy for MemCLR (VS et al., 2023), which was originally tailored for image-based 2D object detection, and extend it to 3D detection scenarios. This involves reading and writing pooled region of interest (RoI) features extracted during the second detection stage and computing the memory-based contrastive loss. We apply all hyperparameters from the original paper by default. Besides, The proposed DPO is a pseudo-labeling-based self-training approach for TTA-3OD. We leverage the self-training paradigm and augmentation strategies from prior works (Yang et al., 2021, 2022; Luo et al., 2021).
Appendix C Quantitative Study
Figure 7. Visualization of box predictions comparing direct inference (No Adapt.), the proposed DPO, and the ground truth labels, across a composite domain shift scenario (Waymo $\rightarrow$ KITTI-C) under heavy snow conditions.
Figure 7 visualizes the box predictions from the source pre-trained 3D detector, the proposed DPO, and the ground truth labels. The detection model, pre-trained on Waymo, is adapted to KITTI-C under conditions simulating heavy snowfall, where many noisy green points are distributed throughout the point clouds. The last row displays images of the same testing scenes with projected 2D ground truth boxes. All detected instances in the point clouds are enclosed in blue 3D boxes. Intuitively, DPO demonstrates its ability to better align with the ground truth labels, evidenced by more accurate locations and fewer false positives. In comparison, direct inference often results in a greater number of boxes that do not contain actual objects, caused by a significant domain shift (i.e., cross-dataset plus heavy snow). Additionally, in the first column, a car obscured behind the white car on the left is missed by the ground truth but detected by both DPO and direct inference. While direct inference achieves high recall, it does so at the cost of numerous false positives (i.e., boxes without actual objects). Conversely, the proposed DPO not only demonstrates high recall but also maintains high precision, effectively reducing false positives and confirming its effectiveness in test-time adaptation for 3D object detection.